### v4

# SynthVLM: Towards High-Quality and Efficient Synthesis of Image-Caption Datasets for Vision-Language Models

Abstract.
Vision-Language Models (VLMs) have recently emerged, demonstrating remarkable vision-understanding capabilities. However, training these models requires large-scale datasets, which brings challenges related to efficiency, effectiveness, quality, and privacy of web data. In this paper, we introduce SynthVLM, a novel data synthesis and curation method for generating image-caption pairs. Unlike traditional methods, where captions are generated from images, SynthVLM utilizes advanced diffusion models and high-quality captions to automatically synthesize and select high-resolution images from text descriptions, thereby creating precisely aligned image-text pairs. To demonstrate the power of SynthVLM, we introduce SynthVLM-100K, a high-quality dataset consisting of 100,000 curated and synthesized image-caption pairs. In both model and human evaluations, SynthVLM-100K outperforms traditional real-world datasets. Leveraging this dataset, we develop a new family of multimodal large language models (MLLMs), SynthVLM-7B and SynthVLM-13B, which achieve state-of-the-art (SOTA) performance on various vision question-answering (VQA) tasks. Notably, our models outperform LLaVA across most metrics with only 18% pretrain data. Furthermore, SynthVLM-7B and SynthVLM-13B attain SOTA performance on the MMLU benchmark, demonstrating that the high-quality SynthVLM-100K dataset preserves language abilities. To facilitate future research, our dataset and the complete data generating and curating methods are open-sourced at https://github.com/starriver030515/SynthVLM.
Synthetic Images, Data Curation, Vision Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD ’25: ACM KDD conference; August 3–7, 2025; Toronto, ON, Canadabooktitle: KDD ’25: ACM KDD conference, August 3–7, 2025, Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
In recent years, with the rapid advancements in large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023) and multimodal large language models (MLLMs) (Zhao et al., 2023; Wu et al., 2023), data management has become a crucial aspect of these technologies (Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023; Huang et al., 2024). At the same time, Bai et al. (2024) also demonstrates that data processing, selection, and management can significantly influence the performance of MLLMs.
Among MLLMs, Vision Large Language Models (VLLMs) achieve competitive performance in traditional multimodal tasks such as image classification (Chen et al., 2024), image understanding (Li et al., 2023b, c), and image captioning (Bai et al., 2023). Most existing VLLMs rely on web images and primarily focus on modifying model architectures to better leverage information from multiple modalities (Li et al., 2023b, c; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024). However, high-quality web data is scarce, and the quality of data significantly impacts the performance of VLLMs. For example, Li et al. (2023d); Wang et al. (2024a) demonstrate that higher-quality training data can enhance the effectiveness of VLLMs. The key to ensuring high-quality data lies in the precise alignment between multimodal inputs, such as the alignment between captions and images.
To address the scarcity of high-quality web data, data synthesis strategies have increasingly been employed, particularly with the advancement of generative models. For instance, Nguyen et al. (2023) utilized BLIP2 to generate numerous image captions, achieving SOTA results on DataComp. In the domain of VLLMs, Chen et al. (2023b) used GPT-4 Vision to produce highly descriptive image captions, resulting in significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLLM performance.
Despite these notable advancements in VLLMs, the challenge of multimodal data persists, particularly when paired image and text are required. Although synthetic captions have been generated for images, the absence of generated images means that the issue of limited high-quality web images remains unresolved. This limitation gives rise to four key challenges that still need to be addressed:
C1. Low Data Quality.
Our evaluations using metrics such as CLIPScore  (Hessel et al., 2021) and training results reveal that existing datasets still align the modalities sub-optimally. Web images often suffer from issues such as blurriness and watermarks. Additionally, methods employing BLIP2  (Nguyen et al., 2023) for generating captions frequently result in logical inconsistencies and unclear grammar, which mislead VLLM training and degrade their language capabilities.
C2.Poor Effectiveness.
Existing datasets frequently lack the necessary diversity and depth required to train VLLMs effectively. For example, many web-scraped datasets consist of low-quality or irrelevant content that fails to capture the complexity of natural language and visual information. Furthermore, the limited scope of current datasets often results in models that struggle to generalize across various contexts.
C3. Low Efficiency.
Methods that rely on manual captioning are both labor-intensive and resource-intensive. Automated solutions such as ShareGPT4v (Chen et al., 2023b), which leverage GPT-4 Vision for labeling, are costly and difficult to scale. Additionally, many existing strategies require the creation of large datasets, such as LLAVA-558K, to boost performance, resulting in significant data redundancy.
C4. Security Risks.
Utilizing web-sourced data introduces numerous security and privacy concerns (Li et al., 2023a; Das et al., 2024). They may contain personal information or copyrighted materials, posing potential legal and ethical challenges. Moreover, the inclusion of sensitive or inappropriate content within training datasets can instigate ethical issues, thereby compromising the models’ integrity and fairness.
To address these challenges, we introduced a new data generation pipeline, SynthVLM. First, we implemented a quality selection process to filter high-quality caption data. Next, we employed advanced diffusion models to generate images based on these captions. For quality assurance, we proposed a new method for evaluating and selecting image-caption pairs by combining CLIPScore (Hessel et al., 2021) and SSIM (Wang et al., 2004), focusing on both image-text alignment and image quality. Our data generation approach achieved superior alignment between images and captions compared to existing methods. With 100K curated synthetic data, we attained SOTA results across multiple benchmarks, utilizing only 18% of the official LLaVA-558K dataset size.
Overall, our contributions are as follows:
  • New Method. We propose a novel two-step image-caption data synthesis method called SynthVLM. Using SynthVLM, we create a high-quality image-caption dataset, SynthVLM-100K. We subsequently use SynthVLM-100K to train both the SynthVLM-7B and SynthVLM-13B models. Our method is both effective and efficient compared to the baseline method. Additionally, our ablation study demonstrates that both steps of SynthVLM are effective.
  • High Data Quality. SynthVLM-100K outperforms baseline datasets in both model and human evaluations. Specifically, SynthVLM-100K achieves superior CLIPScore and SSIMScore, indicating better image-text alignment and image quality. Moreover, it receives better ratings from human annotators, GPT-4, and Intern-VL, further confirming the dataset’s high-quality.
  • Strong Performance on VLLMs. We use only SynthVLM-100K as high-quality pre-training data, outperforms the baseline that uses LLaVA-558K. We achieve SOTA performance on vision-understanding tasks and demonstrate outstanding results on language tasks on MMLU benchmark.
  • Data Privacy. By using synthetic data, we eliminate the need for real personal images, documents, and other sensitive information, thereby ensuring data privacy.
2. Related Work
2.1. Diffusion Model
Denoising diffusion probabilistic models (DDPMs)  (Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023) are a class of generative models renowned for their ability to generate extremely high-quality images. The core idea of DDPMs involves modeling the data distribution by gradually adding Gaussian noise to the input image during the forward process and then predicting and removing this noise to reconstruct the image during the backward process.
Given a source image data distribution $x_{0}\sim q(x_{0})$, Gaussian noise is added over $T$ steps to obtain $x_{T}$. The forward process is defined as:
$q(x_{1},\ldots,x_{T}\mid x_{0}):=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
where $\beta_{t}$ controls the variance of the noise added at each step.
The distribution after $t$ steps can be written as:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
The backward process aims to reconstruct the data by learning a series of Gaussian distributions that approximate the forward process:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are neural networks parameterized by $\theta$.
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022; Su et al., 2023).
Figure 1. The pipeline of the SynthVLM data synthesis method is as follows: First, we filter high-quality image-caption pairs. Next, we synthesize high-quality data and subsequently filter them based on CLIPScore.
2.2. Vision Language Models
The integration of visual knowledge into large language models (LLMs) has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks (Liu et al., 2023d; Zhang et al., 2022; Li et al., 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP (Radford et al., 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters (Liu et al., 2023b, a; Li et al., 2023b, 2022a; Jian et al., 2023; Lu et al., 2023a) are introduced to further integrate different modalities. For example, LLaVA (Liu et al., 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP (Li et al., 2022a, 2023b) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023f) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities. In this work, we propose a new data generation strategy that leverages high-quality caption-image pairs, enabling the model to achieve state-of-the-art (SOTA) results with minimal data. By aligning during the pre-training (PT) phase, we have further enhanced the model’s language capabilities.
2.3. Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data (Touvron et al., 2023; OpenAI, 2023b). In this scenario, the quality and quantity of data become paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data management, cleaning, and annotation (Bai et al., 2024). In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation (Bai et al., 2024). Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection (Bai et al., 2024). For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023c) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023c) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023b) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Figure 2. We compared the SynthVLM-100K dataset with LLaVA-558K. As shown in (a), the generated image can avoid content such as watermarks and advertisements. In (b), the caption mentions a joicy named plate, however in the bottom right image, joicy is actually a book. The generated images better reflect the content of the captions. Additionally, the resolution of the generated images is 1024x1024, which is higher than the existing images, is more beneficial for model training and expansion.
2.4. Data Generation
Data has always been the key driver behind the success of LLMs. Recent advancements of LLMs largely due to the availability of large-scale, diverse, and high-quality datasets for training these models (Lin et al., 2023; Li et al., 2024c, e, 2023g, a, d, b). However, the scarcity of data and the high costs present substantial challenges in obtaining such datasets (Xie et al., 2024; Xu et al., 2023a). Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. Synthetic data holds great potential in building large-scale, high-quality datasets. Researchers have explored multiple approaches, from leveraging differential privacy to creating instruction-tuning frameworks, to enhance the quality, diversity, and utility of synthetic data  (Wang et al., 2024c; Wei et al., 2023; Lou et al., 2023; Yang et al., 2023). A key component in generating high-quality synthetic datasets is precise alignment. Fan et al. (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations. Li et al. (Li et al., 2023e) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1M caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100K high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM-100K dataset with other existing datasets.
3.1. Step1: Synthetic Dataset Construction
In this section, we introduce the image generation pipeline. First, we construct a large pool of captions for selection. Next, we generate high-quality images corresponding to these captions. We then select the best captions from the pool for image-text generation. Utilizing these high-quality captions, we employ diffusion models to generate the images.
Table 1. LCS abbreviates the LAION, CC, and SBU datasets. SynthVLM uses captions to generate images, while others use images to generate captions or manual labeling.
Name Image Source Caption Source Sample
COCO-Caption COCO Human 118K
BLIP-LCS LCS BLIP 558K
ShareGPT4V LCS, COCO, etc GPT4-Vision 100K
ShareGPT4V-PT LCS, COCO, etc Share-Captioner 1246K
SynthVLM Diffusion LCS, COCO, BLIP2-DataComp, etc 1000K
Data Source.
To ensure the diversity of the captions, we combined human-generated and model-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the model-generated captions were predominantly created using the method described in  (Nguyen et al., 2023), which utilizes BLIP2 to regenerate captions for images in the DataComp dataset  (Gadre et al., 2023).
Caption Curation.
To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. The filtering process was performed with ChatGPT, combined with statistical indicators such as N-grams and Perplexity, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions. The formula for calculating CLIPScore is as follows:
(1) ${CLIPScore}(I,C)=\frac{{CLIP}(I)\cdot{CLIP}(C)}{||{CLIP}(I)||\cdot||{CLIP}(C)||}$
where $I$ represents the image, $C$ represents the caption, and ${CLIP}(I)$ and ${CLIP}(C)$ denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and |||||$\cdot$||| | ⋅ | | denotes the norm of the vectors.
We selected the top 40% of image-caption pairs with the highest CLIPScores. These selected captions were included in the candidate caption set. Ultimately, we curated a dataset of 1M captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 1(a).
Image Generation.
After filtering 1M high-quality captions, we employed Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SOTA model capable of efficiently generating high-quality, high-resolution images. We configured SDXL with 60 sampling steps and used 8 A100 GPUs to ensure the superior quality of our synthetic images. Additionally, SynthVLM produces images at a resolution of 1024x1024, effectively addressing the low-resolution issues present in existing datasets. This improvement greatly enhances the quality and utility of the training data across various image generation and recognition tasks.
3.2. Step2: Synthetic Data Selection
In this section, we introduce a novel algorithm for quality control of generated datasets. To better ensure the alignment between images and their corresponding text descriptions, we continue to use CLIPScore. For a given image $I$ and its corresponding text $C$, we first calculate the image-text CLIPScore, ${CLIPScore}(I,C)$, to assess their alignment. Additionally, the generated images have a resolution of $1024\times 1024$, which will be resized to $336\times 336$ for compatibility with CLIP. To account for the potential loss in quality due to this resizing process, we introduce the Structural Similarity Index Measure (SSIM), a metric for image quality control. This will help us better ensure the image quality is preserved after resizing. The SSIM formula is defined as follows:
$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$
Where $x$ and $y$ are the two images being compared, $\mu_{x}$ and $\mu_{y}$ are the mean pixel intensities of the images, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the variances of the images, $\sigma_{xy}$ is the covariance of the images, $C_{1}$ and $C_{2}$ are constants to stabilize the division with weak denominator values.
For a given image $I$ with a resolution of $1024\times 1024$, we first resize it to $336\times 336$, then interpolate to restore it back to $1024\times 1024$. The SSIM value is then computed between the resized image and the original image to quantify the loss introduced by the resizing process. Let $I_{{resized}}$ denote the resized image, where:
$I_{{resized}}={resize}({resize}(I,(336,336)),(1024,1024))$
The SSIMScore is computed as:
$SSIMScore=SSIM(I,I_{{resized}})$
Finally, to combine the CLIPScore and SSIMScore, we apply a weighted sum, with SSIMScore weighted by a factor $\lambda$. The overall score $S$ is given by:
$Weighted\_Score={CLIPScore}(I,C)+\lambda\cdot SSIMScore(I,I_{{resized}})$
In practice, we set $\lambda=0.5$ to balance the contribution of CLIPScore and SSIMScore in the final score.
As shown in Figure 1(b), we initially computed CLIPScores and SSIMScore for the 1M synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset.
Table 2. We employed GPT4-Vision and InternVL to vote on the match between each caption and its corresponding generated image and raw image. The results demonstrate that the generated images align more closely with the captions.
Sample Model Sythetic Image win Web Image win
1K GPT4-Vision 633 367
1K InternVL2 692 308
1K Human 758 242
Table 3. We compared the average CLIPScore, SSIMScore, and their weighted score across our synthetic dataset, ShareGPT4V, COCO-Caption, and BLIP-LCS. The results demonstrate that SynthVLM achieves the highest alignment and image quality.
Datasets # Samples CLIPScore SSIMScore Weighted_Score
COCO-Caption 118K 0.31 0.73 0.67
BLIP-LCS 558K 0.32 0.75 0.70
ShareGPT4V 100K 0.32 0.79 0.71
Synth Dataset 1000K 0.34 0.78 0.73
Curated-Synth Dataset 100K 0.36 0.86 0.79
Table 4. Comparison of data utilization for generating image-caption pairs. This indicates that our SynthVLM have superior efficiency compared to other methods.
Methods SynthVLM LLaVA w/o selection
Dataset Number (k) 100 558 1000
Data Usage 33MB 27GB 330MB
Table 5. Comparison of SynthVLM and LLaVA using the same model structure. We can see SynthVLM outperforms LLaVA on all the evaluation benchmarks.
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
LLaVA-7B Vicuna-1.5-7B 69.3 67.3 30.5 49.9 78.7 62.5 65.3 1484.8 315.6 86.0 36.3
SynthVLM-7B Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
LLaVA-13B Vicuna-1.5-13B 74.2 71.0 35.0 53.6 80.0 63.0 67.7 1531.3 294.5 86.9 52.4
SynthVLM-13B Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
Table 6. Result comparison of MMLU shows that with the synthetic 100k data, our SynthVLM outperforms LLaVA in pure language tasks. This demonstrates the effectiveness of the synthetic data in modality alignment.
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
LLaVA-7B Vicuna-1.5-7B 69.3 36.3 28.6 33.4 39.5 44.5
SynthVLM-7B Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
LLaVA-13B Vicuna-1.5-13B 74.2 52.4 41.9 45.8 62.9 61.8
SynthVLM-13B Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
Figure 3. This is our process and prompt design for match assessment using GPT4-Vision. We consider various aspects, including the quality of the image and the match between the image and the caption, to help GPT4-Vision make a better selection. Based on this process, we compare SynthVLM with existing datasets from the model’s perspective.
3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM-100K dataset. The synthetic data offers high image quality, excellent image-text alignment, superior model ratings, and robust data privacy protection.
High Image Quality
As illustrated in Figure 2, SynthVLM markedly improves image quality by generating images at a resolution of 1024x1024 pixels. This high resolution addresses the prevalent issue of suboptimal image quality in existing datasets, providing high-quality image-caption pairs that are invaluable for training VLLMs. Additionally, SynthVLM-100K effectively eliminates artifacts such as watermarks and advertisements.
Excellent Image-Text Alignment.
As shown in Table 3, the SynthVLM-100K dataset exhibits a higher CLIPScore and SSIMScore compared to existing high-quality web datasets. By selecting curated image-text pairs of superior quality, SynthVLM-100k surpasses datasets such as COCO-Caption, BLIP-LCS, and ShareGPT4V. This demonstrates the outstanding alignment of our dataset.
Excellent Model Rating.
Since our data will be used for VLLMs training, we use VLLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used human annotators, GPT-4 Vision and Intern-VL2 as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 3. The results, presented in Table 2, demonstrate that images generated have better alignment with the caption.
Table 7. The results on the MME benchmark demonstrate that using generated data can still maintain a leading performance in real-world problems, further expanding the application scope of SynthVLM.
Models LLM MME Cognition MME Perception
Reasoning Translation Code Posters Celebrity Scene Landmark Artwork OCR
LLaVA1.5-7B Vicuna-1.5-7B 126.4 57.5 62.5 148.3 132.1 143.0 141.8 123.8 100.0
SynthVLM-7B Vicuna-1.5-7B 143.6 70.0 57.5 157.1 124.2 157.8 156.5 136.2 102.5
LLaVA-13B Vicuna-1.5-13B 119.3 50.0 62.5 155.4 127.4 158.5 165.2 129.3 110.0
SynthVLM-13B Vicuna-1.5-13B 130.0 65.0 62.5 160.4 138.2 162.0 158.8 140.3 104.5
Protect Data Privacy.
A notable advantage of our dataset is its exclusive reliance on captions, which mitigates data privacy concerns. By not using existing images, we ensure that no sensitive or private information associated with those images is compromised. This approach adheres to data privacy best practices, ensuring that our dataset maintains both high quality and ethical integrity.
Through these two training stages, we successfully developed the VLLM model trained by SynthVLM-100K data. Our method is efficient, utilizing only 100K pre-training data while also protecting privacy by leveraging synthetic data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using pure synthetic data.
4. Experiments
In this section, we utilize the image data synthesis system for various tasks. We then aim to answer the following questions to verify the effectiveness, efficiency, and privacy protection of our proposed SynthVLM: Q1: Can our SynthVLM-7B and SynthVLM-13B achieve SOTA performance compared to previous SOTA methods? Q2: How does the efficiency of our SynthVLM compare to previous methods? Q3: Can SynthVLM protect privacy while achieving state-of-the-art performance? Q4: Would using generated data in full impact the application of the model in real-world scenarios? Q5: Do we need the generate module and the quality selection module to enhance model performance?
4.1. Experimental Settings
Models.
We use the LLaVA 1.5 model to validate the effectiveness of our dataset. For the Vision Encoder, we select CLIP 336px, and for the LLM, we use Vicuna 7B and Vicuna 13B. Our training parameters are consistent with LLaVA, and we adopt a two-stage training approach. In the pre-training stage, we train the projector to align the image and text modalities. The SynthVLM-100K dataset described in Section 3.3 is used for this purpose. In the SFT stage, we further train the projector alongside the LLM to improve visual understanding capabilities.
Datasets.
For LLaVA-BaseLine-7B and LLaVA-BaseLine-13B, we use the LLaVA 558k dataset for pre-training and the LLaVA 665k dataset for SFT. For SynthVLM-7B and SynthVLM-13B, we use the SynthVLM-100K dataset for pre-training and the LLaVA 665k dataset for SFT.
Benchmarks.
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQAI, MMVet, VizWiz, VQAv2, GQA, MME, and PoPE for a comprehensive evaluation. For language benchmarks, we select MMLU and SQA to assess language understanding abilities.
Settings.
We primarily adhered to the hyperparameters specified in the official repository for the LLaVA model evaluations. All experiments were conducted on an 8x NVIDIA A100 GPU machine equipped with a 120-core CPU and 960GB of memory.
4.2. Synthetic Data Achieves SOTA Performance
To address Q1, we trained the model described in experimental settings. From Table 5, it is evident that SynthVLM model outperforms the Baseline across all evaluation benchmarks on both 7B and 13B. Specifically, SynthVLM model achieves SOTA results on visual benchmarks such as SQAI, MMVet, VizWiz, VQAv2, GQA, MME, MMB, and PoPE. Furthermore, SynthVLM model also excels in pure language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
These results demonstrate the strong alignment capability of our synthetic data. Additionally, this provides a new paradigm for effective visual understanding model modality alignment using generated data. During pre-training, it is common to train on all available data due to uncertainty about data selection. Here, we offer 100k high-quality synthetic data as a benchmark for selecting aligned generated data efficiently.
Figure 4. From (a), it is evident that synthetic images can avoid displaying real license plates and ticket information. In contrast, (b) contains actual license plates and ticket information, which can potentially lead to privacy issues.
4.3. Efficient Vision Language Alignment
To address Q2, we examine the computational resource usage during training and evaluate the data utilization efficiency for generating image-caption pairs.
Computational Resource Usage
By integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SOTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
Data Utilization Efficiency
Table 4 compares the data utilization of our proposed method with the conventional BLIP-LCS method, which employs the BLIP model for caption generation. Our approach is significantly more efficient, requiring only 330 MB for captions to generate 1,000,000 image-caption pairs, compared to the traditional methods, which may exceed 50 GB for images. This substantial difference highlights the efficiency of relying solely on captions rather than images in generating image-caption pairs.
Table 8. Ablation study of visual understanding ability and pure language ability. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
SynthVLM-7B Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 67.0$\downarrow$ 31.2$\downarrow$ 46.8$\downarrow$ 79.3$\downarrow$ 62.9$\downarrow$ 66.2$\downarrow$ 1488.8$\downarrow$ 327.5$\downarrow$ 86.2$\downarrow$ 39.1$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
SynthVLM-13B Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
w/o generation Vicuna-1.5-13B 73.6$\downarrow$ 71.4$\downarrow$ 33.0$\downarrow$ 53.6$\downarrow$ 80.0 63.4$\downarrow$ 67.5$\downarrow$ 1514.3$\downarrow$ 295.7$\downarrow$ 88.2$\downarrow$ 53.6$\downarrow$
w/o selection Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
Table 9. Ablation study of modality alignment. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
SynthVLM-7B Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 39.1$\downarrow$ 30.0$\downarrow$ 36.6$\downarrow$ 43.1$\downarrow$ 47.3$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 40.6$\downarrow$ 30.8$\downarrow$ 37.2$\downarrow$ 45.3$\downarrow$ 48.9$\downarrow$
SynthVLM-13B Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
w/o generation Vicuna-1.5-13B 74.1$\downarrow$ 53.6$\downarrow$ 43.5$\downarrow$ 48.2$\downarrow$ 63.1$\downarrow$ 61.8$\downarrow$
w/o selection Vicuna-1.5-13B 73.6$\downarrow$ 53.0$\downarrow$ 42.9$\downarrow$ 46.8$\downarrow$ 63.8$\downarrow$ 61.3$\downarrow$
4.4. Privacy Protection Synthetic Data
To address Q3, we compare the synthetic image and the original image in Figure 4. Synthetic data offers significant advantages in protecting data privacy. In Figure 4, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
4.5. Capability to Solve Real-World Problems
A critical issue when using generated data is whether the model loses its ability to solve real-world problems. To address Q4, we utilized the MME Benchmark to assess the model’s performance in various real-world scenarios, such as artwork, celebrity, code reasoning, landmarks, and posters. The specific results are presented in Table 8. Using SynthVLM-100K, our model performs consistently well, matching or even surpassing the baseline across most real-world tasks. We attribute this success to the diversity of the captions and the generalization ability of the Diffusion model.
4.6. Ablation Study
To address Q5, we conducted the following ablation study. Specifically, we conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline. In this section, we controlled the number of samples in each experimental group to be 100K.
Excluding Data Generation Module
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 8 and 9, labeled as ”w/o generation”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
Excluding Data Selection Module
The absence of the data selection module similarly leads to a noticeable decline in performance, indicated as ”w/o selection” in Tables 8 and 9. Given the inherent randomness of diffusion models, which inevitably generate some low-quality images, the data selection module is crucial for removing these subpar elements.
Overall, the ablation study highlights the critical role of the data generation and the data selection modules in SynthVLM. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the SynthVLM model.
5. Conclusion
Data is crucial for training VLLMs. However, high-quality web data is limited. In this paper, we introduce a novel image synthesis pipeline to generate pre-training data for VLLMs. Our synthetic images are of higher quality as they are free from watermarks and advertisements. Furthermore, our synthetic image-caption pairs exhibit superior CLIPScore and SSIMScore, reflecting better image-caption alignment and image quality. Remarkably, the SynthVLM model, trained with only 18% synthetic data, surpasses the LLaVA baseline model trained on the full dataset. Moreover, it efficiently achieves SOTA alignment capabilities while preserving the language proficiency of VLLMs. By employing synthetic datasets, SynthVLM also safeguards data privacy, underscoring the practicality and effectiveness of our data synthesis pipeline.